Angle "P" measures 30 degrees more than its supplement. What is the measure of angle "P" in dregrees?
Answer by mananth(16946)

let angle p = x deg
complement = 90-x
..
x= 90-x +30
x+x=120
2x=120
/2
x= 60 deg.
...
m.ananth@hotmail.ca